1.   A survey is conducted on attitudes towards drinking. A random sample of eight married couples is selected, and the husbands and wives respond to an attitude-toward-drinking scale. (20 Points)The scores are as follows:

HUSBAND SCORE - 16 , 20, 10, 15, 8, 19, 14, 15

WIFE SCORE - 15, 18, 13, 10, 12, 16, 11, 12


1
Expert's answer
2021-07-19T05:45:57-0400
"\bar{x}_1=\dfrac{1}{8}(16+ 20+10+15+8+19+14+15)"

"=14.625"

"s_1^2=\dfrac{1}{8-1}((16-14.625)^2+ (20-14.625)^2"

"+(10-14.625)^2+(15-14.625)^2+(8-14.625)^2"

"+(19-14.625)^2+(14-14.625)^2+(15-14.625)^2)"

"=\dfrac{115.875}{7}\approx16.5535"

"s_1=\sqrt{s_1^2}\approx4.0686"



"\bar{x}_2=\dfrac{1}{8}(15+ 18+13+10+12+16+11+12)"

"=13.375"

"s_2^2=\dfrac{1}{8-1}((15-13.375)^2+ (18-13.375)^2"

"+(13-13.375)^2+(10-13.375)^2+(12-13.375)^2"

"+(16-13.375)^2+(11-14.625)13.375^2+(12-13.375)^2)"

"=\dfrac{51.875}{7}\approx7.4107"

"s_2=\sqrt{s_2^2}\approx2.7223"


A F-test is used to test for the equality of variances.

The critical values for a significance level of "\alpha=0.05," and degrees of freedom "df_1=7"

and "df_2=7" are "F_U=F_{1-\alpha/2; df_1, df_2}=4.995" and "F_L=F_{\alpha/2; df_1, df_2}=0.2."

The F-statistic is calculated by dividing the sample variances, as shown below:


"F=\dfrac{s_1^2}{s_2^2}=\dfrac{4.0686}{2.7233}\approx2.2337"

Since "F_L=0.2<F=2.2337<F_U=4.995," then we fail to reject the null hypothesis of equal population variances, and hence, we assume that the population variances are equal. 


1. We need to construct the 95% confidence interval for the difference between the population means "\mu_1-\mu_2," for the case that the population standard deviations are not known. 

We assume that the population variances are equal. Therefore, the corresponding degrees of freedom are computed as follows:


"df=df_1+df_2=7+7=14"

The critical value for "\alpha=0.05" and "df=14" is "t_c=2.144787."

The corresponding confidence interval is computed as shown below:

Since the population variances are assumed to be equal, we need to compute the pooled standard deviation, as follows:


"s_p=\sqrt{\dfrac{(n_1-1)s_1^2+(n_2-1)s_2^2}{n_1+n_2-2}}"

"=\sqrt{\dfrac{(8-1)\dfrac{115.875}{7}+(8-1)\dfrac{51.875}{7}}{8+8-2}}\approx3.4615"

Since we assume that the population variances are equal, the standard error is computed as follows:


"SE=s_p\sqrt{\dfrac{1}{n_1}+\dfrac{1}{n_2}}=3.4615\sqrt{\dfrac{1}{8}+\dfrac{1}{8}}\approx1.7308"

Now, we finally compute the confidence interval:


"CI=(\bar{x_1}-\bar{x_2}-t_c\times SE, \bar{x_1}-\bar{x_2}+t_c\times SE)"

"=(14.625-13.375-2.144787\times 1.7308,"

"14.625-13.375+2.144787\times 1.7308)"

"=(-2.4622,4.9622)"

Therefore, based on the data provided, the 95% confidence interval for the difference between the population means "\mu_1-\mu_2" is "-2.4622<\mu_1-\mu_2<4.9622," which indicates that we are 95% confident that the true difference between population means is contained by the interval "(-2.4622,4.9622)."


2. The following null and alternative hypotheses need to be tested:

"H_0:\mu_1=\mu_2"

"H_1:\mu_1\not=\mu_2"

The rejection region for this two-tailed test is "R=\{t:|t|>2.144787\}."

Since it is assumed that the population variances are equal, the t-statistic is computed as follows:


"t=\dfrac{\bar{x}_1-\bar{x}_2}{\sqrt{\dfrac{(n_1-1)s_1^2+(n_2-1)s_2^2}{n_1+n_2-2}(\dfrac{1}{n_1}+\dfrac{1}{n_2})}}=0.722"

Since it is observed that "|t|=0.722<2.144787=t_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach: The p-value for two-tailed, "\alpha=0.05, df=14, t=0.722" is "p=0.482181," and since "p=0.482181>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu_1" is different than "\mu_2," at the "\alpha=0.05" significance level.